OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -3, an = 2 ● an-1

myininaya (myininaya):

replace n with 2

myininaya (myininaya):

\[a_1=-3 \\ a_n=2 \cdot a_{n-1} \\ \text{ replacing } n \text{ with } 2 \\ a_2=2 \cdot a_{2-1} \\ a_2=2 \cdot a_1 \text {since 2-1=1} \\ \text{ now recall } a_1=-3 \\ \text{ replace } a_1 \text{ with -3 } \\ a_2=2 (-3)=-6 \text{, -6 is the second term }\]

myininaya (myininaya):

try finding the 3rd term by replacing n with 3 instead of 2

OpenStudy (anonymous):

-12 3 -24 4 -48 5 -96 6

myininaya (myininaya):

checking... \[a_3=2 \cdot a_{3-1} \\ a_3=2 \cdot a_2 \\ \text{ but remember } a_2=-6 \\ a_3=2(-6)=-12\] your third term is right anyways we are just multiplying previous term by 2 so we started with -3 -3(2)=-6 -6(2)=-12 -12(2)=-24 -24(2)=-48 -48(2)=-96 yep seems great

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I might have a few more questions later

myininaya (myininaya):

hey on the previous question where you asked what 9+10 like was that really the question?

OpenStudy (anonymous):

lol, no @myininaya

myininaya (myininaya):

Like if you can give me the actual question I might be able to help you with it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Im just studying for my semester exam so it can be any type of question. I dont have any right now tho

myininaya (myininaya):

oh ok

OpenStudy (anonymous):

actually I lied, I'll post a new question